The common stock of Dayton Repair sells for $47.92 a share. The stock is expected to pay $2.28 per share next year when the annual dividend is distributed. The company increases its dividends by 1.65 percent annually. What is the market rate of return on this stock?
Answer #1

Rate of return = ($2.28 / $47.92) + 0.0165

Rate of return = 0.0641 or 6.41%
Answer #2

Given :


P0 = 47.92($)

D1 = 2.28 ($)

g = 1.65 % = 0.0165



Now,


Dividend yield = D1/P0 

= 2.28/47.92

= 0.0476


Market rate of return, r 

= Dividend yield + g

= 0.0476+ 0.0165

= 0.0641

= 6.41% 


So, market rate of return = 6.41 % (ANSWER).


answered by: Tulsiram Garg